OpenStudy (jessiegonzales):

MEDAL AND FANN!!!!! 20. What is an equation for the line that passes through points (-1,-4) and (1,4)? Write the equation in slope-intercept form.

OpenStudy (kagıtucak):

The general formula of a line is y=mx+n Now, lets put the points so that they satisfy this equation For (-1,-4): -4=-m+n call this (1) For (1,4): 4=m+n call this (2) Now lets add (1) and (2) we get; 0=2n then n=0 Now put n=0 into (1) -4=-m then m=4 Thus our line's equation is y=4x

OpenStudy (jessiegonzales):

awesome is that all?

OpenStudy (kagıtucak):

Yes, that's it.

OpenStudy (jessiegonzales):

thanks so much are the steps in order?

OpenStudy (kagıtucak):

Yes, they are in order you can use them.